# Introducing Routing Functions to Vision-Language Parameter-Efficient Fine-Tuning with Low-Rank Bottlenecks

### v1

@@ Appendix 0.B Additional Results with RoBERTa/GPT2 + ViT
$x_{t}\circ x^{\prime}_{v}$ $8.724.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.7}}$ $4.6\underline{23.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.6}}}$ $9.446.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.4}}$ $22.984.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+22.9}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.1}}}$ $9.7439.16_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.74}}$ $2.022.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.0}}$ $2.621.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.6}}$ $42.8_{+0.1}$ $13.375.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+13.3}}$ $3.716.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.7}}$ $3.3435.70_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.34}}$
$x_{t}+x^{\prime}_{v}$ $5.121.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.1}}$ $3.021.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.0}}$ $5.442.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.4}}$ $14.475.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+14.4}}$ $1.916.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.9}}$ $5.7635.18_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.76}}$ $5.225.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.2}}$ $4.2\underline{23.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.2}}}$ $4.9\underline{47.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.9}}}$ $26.1\underline{88.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+26.1}}}$ $4.7\underline{17.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.7}}}$ $9.02\underline{40.38_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.02}}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $10.2\underline{26.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.2}}}$ $4.7\mathbf{23.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.7}}}$ $9.9\mathbf{46.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.9}}}$ $23.8\underline{85.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+23.8}}}$ $3.2\mathbf{17.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.2}}}$ $10.36\underline{39.78_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.36}}}$ $5.3\underline{25.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.3}}}$ $3.923.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.9}}$ $3.546.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.5}}$ $20.782.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+20.7}}$ $4.317.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.3}}$ $7.5438.90_{\color[rgb]{0.1914,0.3359375,0.9414}{+7.54}}$
$x_{t}x^{\prime\prime}_{v}$ $10.7\mathbf{26.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.7}}}$ $4.523.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.5}}$ $9.8\underline{46.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.8}}}$ $24.2\mathbf{85.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+24.2}}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.1}}}$ $10.46\mathbf{39.88_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.46}}}$ $6.2\mathbf{26.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.2}}}$ $4.9\mathbf{24.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.9}}}$ $6.0\mathbf{48.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.0}}}$ $27.9\mathbf{90.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+27.9}}}$ $5.0\mathbf{17.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.0}}}$ $10.00\mathbf{41.36_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.00}}}$
@@
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 3.932M $2.030.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.0}}$ $0.825.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+0.8}}$ $1.453.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.4}}$ $7.299.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+7.2}}$ $1.519.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.5}}$ $2.6645.68_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.66}}$
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 4.746M $1.330.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.3}}$ $0.825.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+0.8}}$ $1.153.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.1}}$ $6.899.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.8}}$ $1.419.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.4}}$ $2.3645.38_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.36}}$
@@
$\text{Adapter}^{\dagger}$ $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 1.830M $0.130.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+0.1}}$ 25.6 53.3 $98.8_{-1.0}$ $18.9_{-0.2}$ $45.48_{-0.22}$
@@
As presented in Figure B8, with the addition of routing function, the changes in down-project mapping $W_{down}$ are more drastic than changes in up-project mapping $W_{up}$. It shows that the routing functions have bigger impact on $W_{down}$ as training progresses, which is desired as we are learning VL alignment in the low-rank bottleneck. This intriguing phenomenon reveals properties of the learning dynamics of routing functions.
@@ Appendix 0.C Additional Results with CLIP-BART
In this section we report the standard deviations obtained by routing function ($x_{t}(x_{v})^{\mathsf{T}}x_{v}$) and VL-Adapter (as reported in Hu_2023_ICCV ). For the detailed evaluation on routing functions with single Adapter/LoRA for CLIP-BART, we refer to Sec. 5.3 (Tab. 7).
Variations in CLIP-BART experiments To verify whether the results produced by the routing functions are robust enough, we are also interested in the variations over different runs. Limited to resources, we select one routing function, that is, $x_{t}(x_{v})^{\mathsf{T}}x_{v}$, and report the mean and standard deviation over three runs. We use the same seeds as adopted in Hu_2023_ICCV , and report the results in Tab. C15.
@@
Functions VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg. VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg.
$\text{None}^{\dagger}$ $65.15\pm 0.16$ $53.66\pm 0.84$ $72.58\pm 0.73$ $115.01\pm 0.26$ $76.60\pm 0.32$ $65.76\pm 0.28$ $54.16\pm 0.44$ $73.19\pm 0.71$ $114.61\pm 0.26$ $76.93\pm 0.25$
$x_{v}(x_{v})^{\mathsf{T}}x_{v}$ $64.94\pm 0.09$ $53.56\pm 0.31$ $73.60\pm 0.36$ $117.80\pm 0.34$ $77.48\pm 0.11$ $65.84\pm 0.32$ $53.65\pm 0.47$ $74.31\pm 0.09$ $117.65\pm 0.39$ $77.87\pm 0.18$
As presented in Tab. C15, the results obtained by the routing function are consistent. We even obtain much lower standard deviations in many tasks (e.g. on GQA&$\text{NLVR}^{2}$ for Single LoRA, on $\text{NLVR}^{2}$ for Single Adapter). We also obtain lower standard deviations on average scores for both Single LoRA and Single Adapter. The results show that our method is robust with low variations in performance.

### v2

@@ Appendix 0.B Additional Results with RoBERTa/GPT2 + ViT
$x_{t}\circ x^{\prime}_{v}$ $8.724.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.7}}$ $4.6\underline{23.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.6}}}$ $9.446.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.4}}$ $22.984.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+22.9}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.1}}}$ $9.7439.16_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.74}}$ $2.022.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.0}}$ $2.621.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.6}}$ $42.8_{+0.1}$ $13.375.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+13.3}}$ $3.716.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.7}}$ $3.3435.70_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.34}}$
$x_{t}+x^{\prime}_{v}$ $5.121.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.1}}$ $3.021.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.0}}$ $5.442.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.4}}$ $14.475.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+14.4}}$ $1.916.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.9}}$ $5.7635.18_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.76}}$ $5.225.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.2}}$ $4.2\underline{23.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.2}}}$ $4.9\underline{47.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.9}}}$ $26.1\underline{88.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+26.1}}}$ $4.7\underline{17.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.7}}}$ $9.02\underline{40.38_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.02}}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $10.2\underline{26.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.2}}}$ $4.7\mathbf{23.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.7}}}$ $9.9\mathbf{46.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.9}}}$ $23.8\underline{85.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+23.8}}}$ $3.2\mathbf{17.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.2}}}$ $10.36\underline{39.78_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.36}}}$ $5.3\underline{25.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.3}}}$ $3.923.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.9}}$ $3.546.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.5}}$ $20.782.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+20.7}}$ $4.317.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.3}}$ $7.5438.90_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+7.54}}$
$x_{t}x^{\prime\prime}_{v}$ $10.7\mathbf{26.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.7}}}$ $4.523.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.5}}$ $9.8\underline{46.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.8}}}$ $24.2\mathbf{85.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+24.2}}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.1}}}$ $10.46\mathbf{39.88_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.46}}}$ $6.2\mathbf{26.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.2}}}$ $4.9\mathbf{24.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.9}}}$ $6.0\mathbf{48.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.0}}}$ $27.9\mathbf{90.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+27.9}}}$ $5.0\mathbf{17.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.0}}}$ $10.00\mathbf{41.36_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.00}}}$
@@
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 3.932M $2.030.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.0}}$ $0.825.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+0.8}}$ $1.453.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.4}}$ $7.299.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+7.2}}$ $1.519.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.5}}$ $2.6645.68_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.66}}$
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 4.746M $1.330.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.3}}$ $0.825.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+0.8}}$ $1.153.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.1}}$ $6.899.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.8}}$ $1.419.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.4}}$ $2.3645.38_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.36}}$
@@
$\text{Adapter}^{\dagger}$ $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 1.830M $0.130.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+0.1}}$ 25.6 53.3 $98.8_{-1.0}$ $18.9_{-0.2}$ $45.48_{-0.22}$
Routing functions in the PEFT modules outperform cross-attention when the number of parameters is similar. Taking, e.g., $x_{t}(x_{v})^{T}x_{v}$ in Tab. 5, the computation is: $W_{down}x_{H}(W_{down}x_{R})^{T}W_{down}x_{R}$; while for cross-attention, we compute as $\phi(W^{Q}_{down}x_{H}(W^{K}_{down}x_{R})^{T})W^{V}_{down}x_{R}$, where $\phi(\cdot)$ refers to a scaling function in cross-attention and $Q$, $K$ and $V$ refer to the query, key and value matrices, respectively. As discovered by TextSpan (Gandelsman et al. ICLR2024), the key to maintain the performance of attention lies in the number of dimensions of the embedding space where they define a lower bound of embedding dimensions to successfully capture the information in the attention heads. We denote the dimension of the low-rank bottleneck as $r_{W_{down}}$. Let the dimension of the model be $d$. For cross attention, the number of parameters is $\sum_{i\in\{Q,K,V\}}d\times r_{W^{i}_{down}}+r_{W^{i}_{down}}$, with the last term referring to the bias term. For routing functions, the number of parameters is $d\times r_{W_{down}}+r_{W_{down}}$. Thus, for the same number of parameters, routing functions allow for a larger dimension of the bottleneck (for the example above, roughly by a factor of 3, e.g. $r_{W^{i}_{down}}\approx r_{W_{down}}/3$). Alternatively, the same bottleneck dimension can be reached with fewer parameters, reducing the risk of overfitting.
@@
As presented in Fig. B8, with the addition of routing function, the changes in down-project mapping $W_{down}$ are more drastic than changes in up-project mapping $W_{up}$. It shows that the routing functions have bigger impact on $W_{down}$ as training progresses, which is desired as we are learning VL alignment in the low-rank bottleneck. This intriguing phenomenon reveals properties of the learning dynamics of routing functions.
@@ Appendix 0.C Additional Results with CLIP-BART
In this section we report the standard deviations obtained by routing function ($x_{t}(x_{v})^{\mathsf{T}}x_{v}$) and VL-Adapter (as reported in [11]). For the detailed evaluation on routing functions with single Adapter/LoRA for CLIP-BART, we refer to Sec. 5.3 (Tab. 7).
Variations in CLIP-BART experiments To verify whether the results produced by the routing functions are robust enough, we are also interested in the variations over different runs. Following [11], we report the mean and standard deviation over three runs. The results from the most performant routing functions are presented in Tab. C15. The improvements in all settings are statistically significant, except for GQA w/ single Adapter. Large gains are seen for $\text{NLVR}^{2}$ and COCO Cap.
Table C15: Significance test for best routing functions ($\text{mean}_{\pm\text{std. dev.}}$). We mark the significantly better results in bold. We report the average accuracy of VQA, GQA & $\text{NLVR}^{2}$ as Avg. dis. For full results please refer to Tab. C16 and Tab. C17.
VQA GQA $\text{NLVR}^{2}$ Avg. dis. COCO Cap.
Single LoRA None[11] $65.15_{\pm 0.16}$ $53.66_{\pm 0.84}$ $72.58_{\pm 0.73}$ 63.80 $115.01_{\pm 0.26}$
routing $\mathbf{65.68_{\pm 0.01}}$ $\mathbf{53.96_{\pm 0.12}}$ $\mathbf{73.60_{\pm 0.36}}$ $\mathbf{64.35_{\pm 0.22}}$ $\mathbf{119.26_{\pm 0.25}}$
function ($x_{t}\circ x^{\prime}_{v}$) ($x_{t}\circ x^{\prime}_{v}$) ($x_{t}(x_{v})^{\mathsf{T}}x_{v}$) ($x_{t}\circ x^{\prime}_{v}$) ($x_{t}x^{\prime\prime}_{v}$)
Adpt. None[11] $65.76_{\pm 0.28}$ $54.16_{\pm 0.44}$ $73.19_{\pm 0.71}$ 64.37 $114.61_{\pm 0.26}$
routing $\mathbf{65.92_{\pm 0.20}}$ $54.54_{\pm 0.24}$ $\mathbf{74.52_{\pm 0.36}}$ $\mathbf{64.83_{\pm 0.05}}$ $\mathbf{118.86_{\pm 0.15}}$
function ($x_{t}\circ x^{\prime}_{v}$) ($x_{t}\circ x^{\prime}_{v}$) ($x_{t}\circ x^{\prime}_{v}$) ($x_{t}\circ x^{\prime}_{v}$) ($x_{t}x^{\prime\prime}_{v}$)
Multi. Adpt. None[37] 65.4 54.0 69.8 63.0 114.3
routing $\mathbf{66.39_{\pm 0.04}}$ $\mathbf{54.22_{\pm 0.19}}$ $\mathbf{71.83_{\pm 0.80}}$ $\mathbf{63.96_{\pm 0.26}}$ $\mathbf{117.81_{\pm 0.75}}$
function (comb1) (comb1) (comb3) (comb3) (comb1)
@@
Functions VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg. Avg. dis. VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg. Avg. dis.
$\text{None}^{\dagger}$ $65.15_{\pm 0.16}$ $53.66_{\pm 0.84}$ $72.58_{\pm 0.73}$ $115.01_{\pm 0.26}$ $76.60_{\pm 0.32}$ 63.80 $65.76_{\pm 0.28}$ $54.16_{\pm 0.44}$ $73.19_{\pm 0.71}$ $114.61_{\pm 0.26}$ $76.93_{\pm 0.25}$ 64.37
$x_{t}\circ x^{\prime}_{v}$ $65.68_{\pm 0.01}$ $53.96_{\pm 0.12}$ $73.42_{\pm 0.63}$ $113.94_{\pm 0.43}$ $76.75_{\pm 0.06}$ $64.35_{\pm 0.22}$ $65.92_{\pm 0.20}$ $54.34_{\pm 0.17}$ $74.23_{\pm 0.23}$ $114.38_{\pm 0.60}$ $77.22_{\pm 0.16}$ $64.83_{\pm 0.05}$
$x_{t}+x^{\prime}_{v}$ $65.14_{\pm 0.10}$ $53.73_{\pm 0.26}$ $73.51_{\pm 0.22}$ $114.96_{\pm 0.53}$ $76.84_{\pm 0.07}$ $64.13_{\pm 0.13}$ $65.89_{\pm 0.10}$ $54.18_{\pm 0.53}$ $73.90_{\pm 0.28}$ $114.39_{\pm 0.77}$ $77.09_{\pm 0.35}$ $64.66_{\pm 0.21}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $64.94_{\pm 0.09}$ $53.56_{\pm 0.31}$ $73.60_{\pm 0.36}$ $117.80_{\pm 0.34}$ $77.48_{\pm 0.11}$ $64.03_{\pm 0.14}$ $65.84_{\pm 0.32}$ $53.65_{\pm 0.47}$ $74.31_{\pm 0.09}$ $117.65_{\pm 0.39}$ $77.87_{\pm 0.18}$ $64.03_{\pm 0.14}$
$x_{t}x^{\prime\prime}_{v}$ $64.84_{\pm 0.06}$ $53.13_{\pm 0.37}$ $72.98_{\pm 0.44}$ $119.26_{\pm 0.25}$ $77.55_{\pm 0.10}$ $63.65_{\pm 0.06}$ $65.83_{\pm 0.22}$ $53.61_{\pm 0.33}$ $73.27_{\pm 1.24}$ $118.86_{\pm 0.15}$ $77.90_{\pm 0.25}$ $64.24_{\pm 0.38}$
[…]
Routing Routing Functions Performance
Functions VQA GQA $\text{NLVR}^{2}$ COCO Cap. VQA GQA $\text{NLVR}^{2}$ COCO Cap. Avg. Avg. dis.
$\text{None}^{\ddagger}$ 65.4 54.0 69.8 114.3 75.9 63.07
comb1 $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $x_{t}x^{\prime\prime}_{v}$ $66.39_{\pm 0.04}$ $54.11_{\pm 0.37}$ $69.97_{\pm 0.74}$ $117.81_{\pm 0.75}$ $77.07_{\pm 0.37}$ $63.49_{\pm 0.31}$
comb2 $x_{t}+x^{\prime}_{v}$ $x_{t}+x^{\prime}_{v}$ $x_{t}(x_{v})^{T}x_{v}$ $x_{t}x^{\prime\prime}_{v}$ $65.04_{\pm 0.20}$ $53.00_{\pm 0.40}$ $70.96_{\pm 0.41}$ $117.00_{\pm 0.51}$ $76.50_{\pm 0.21}$ $63.07_{\pm 0.11}$
comb3 $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}x^{\prime\prime}_{v}$ $66.28_{\pm 0.20}$ $53.77_{\pm 0.45}$ $71.83_{\pm 0.80}$ $117.66_{\pm 0.67}$ $77.46_{\pm 0.35}$ $63.96_{\pm 0.26}$
comb4 $x_{t}+x^{\prime}_{v}$ $x_{t}+x^{\prime}_{v}$ $x_{t}\circ x^{\prime}_{v}$ $x_{t}x^{\prime\prime}_{v}$ $65.32_{\pm 0.13}$ $53.15_{\pm 0.15}$ $71.00_{\pm 0.11}$ $117.40_{\pm 0.93}$ $76.72_{\pm 0.21}$ $63.16_{\pm 0.07}$
As presented in Tab. C16 and Tab. C17, the results obtained by the routing function are consistent. We even obtain much lower standard deviations in many tasks (e.g. on GQA&$\text{NLVR}^{2}$ for Single LoRA, on $\text{NLVR}^{2}$ for Single Adapter). We also obtain lower standard deviations on average scores for both Single LoRA and Single Adapter. The results show that our method is robust with low variations in performance.